OpenStudy (anonymous):

Find all numbers c that satisfy the conclusion of Rolle's Theorem for the following function and interval f(x)= 2sin(2pi(x)), [-1,1]

OpenStudy (anonymous):

take the derivative, get \[4\pi \cos(\pi x)\] set it equal to zero and solve